### v2

# Can GPT models Follow Human Summarization Guidelines? A Study for Targeted Communication Goals

Abstract
This study investigates the ability of GPT models (ChatGPT, GPT-4 and GPT-4o) to generate dialogue summaries that adhere to human guidelines. Our evaluation involved experimenting with various prompts to guide the models in complying with guidelines on two datasets: DialogSum (English social conversations) and DECODA (French call center interactions). Human evaluation, based on summarization guidelines, served as the primary assessment method, complemented by extensive quantitative and qualitative analyses. Our findings reveal a preference for GPT-generated summaries over those from task-specific pre-trained models and reference summaries, highlighting GPT models’ ability to follow human guidelines despite occasionally producing longer outputs and exhibiting divergent lexical and structural alignment with references. The discrepancy between ROUGE, BERTScore, and human evaluation underscores the need for more reliable automatic evaluation metrics.111The generated summaries and human annotations will be made publicly available at https://github.com/yongxin2020/LLM-Sum-Guidelines.
Keywords Large Language Models Human Summarization Guidelines Dialogue Summarization
1 Introduction
Although instruction-tuned Large Language Models (LLMs) have shown impressive performance in several benchmark datasets, they still struggle with various challenging tasks [1, 2, 3].
For instance, general-purpose models like ChatGPT generally underperform in summarization tasks compared to task-specific pre-trained models in automatic metrics [3, 4, 5]. In dialogue sumarization, ChatGPT demonstrates poorer performance than state-of-the-art (SOTA) fine-tuned models on the ROUGE metric for datasets like SAMSum [6] and DialogSum [7].
However, the underlying causes of GPT models’ underperformance on automatic summarization metrics remain poorly understood. Human summarization is a complex cognitive process that involves understanding, condensing, and conveying the essential information of a source text. During the data annotation phase, human approach summarization by adhering to the provided guidelines. Given that human annotators have to follow these guidelines when creating reference summaries, we speculate that the poor performance of LLMs could be attributed to the lack of explicit summarization instructions tailored to specific communication goals. This is particularly critical for dialogue summarization, where targeted communication goals play a significant role.
Even though dialogue summarization is a well-established task, it remains a challenging task due to the absence of a consensus for what constitutes an ideal dialogue summary [8]. The approach to dialogue summarization is heavily influenced by specific communication objectives, which vary across different contexts such as meetings, or customer service interactions. A review of guidelines from major dialogue summarization datasets reveals that each corpus defines its own set of objectives for creating reference summaries, applying different summary criteria to meet specific needs [9]. For instance, in the context of customer service, while TWEETSUMM [10] offers both extractive and abstractive summaries, CSDS [11] provides three distinct summaries for each dialogue: an overall summary and two role-oriented summaries (user and agent).
Our key contributions in evaluating GPT models’ adherence to human summarization guidelines are as follows:
Developing a range of prompts, from simplified instructions to detailed human-annotator guidelines, for summarization tasks;
Evaluating three GPT models against task-specific pre-trained models on two multilingual datasets with distinct communication goals (DialogSum and DECODA);
Assessing performance using automatic metrics (ROUGE, BERTScore, and LLM-as-judge) and a task-aligned human evaluation framework;
Performing comprehensive analyses to assess and explain model performance, identifying both capabilities and limitations in meeting targeted communication objectives.
2 Experimental Setup
Datasets.
We used two datasets covering different communication goals, languages and contexts: DialogSum [7] and DECODA [12].
DialogSum consists of social conversations in English and includes 12,460, 1,500 and 500 samples in its training, validation and test splits, respectively. It was recently used in a challenge [13], which favors comparison of results.
DECODA [12] is a call-center human-human spoken conversation corpus in French, collected from the RATP (Paris public transport authority) and mostly deals with customer inquiries and agent responses. The corpus was proposed as a pilot task at Multiling 2015 [12]. The test set used in this study contains 100 conversations with 212 synopses, i.e. references.
Models and Parameters.
We used OpenAI ChatGPT (gpt-3.5-turbo)222OpenAI GPT-3.5: https://platform.openai.com/docs/models/gpt-3-5, the latest version available used in September 2023 is gpt-3.5-turbo-0613, GPT-4333OpenAI GPT-4: https://platform.openai.com/docs/models/gpt-4, version gpt-4-0613, and GPT-4o444gpt-4o-2024-08-06: https://platform.openai.com/docs/models#gpt-4o APIs for our experiments.
gpt-3.5-turbo, the most capable GPT-3.5 variant, is optimized for conversational tasks while maintaining strong performance on standard completion tasks. The model handles inputs up to tokens, compared to GPT-4’s -token limit and GPT-4o’s -token context windows. To ensure a stable output, we configured the temperature parameter at while maintaining the default values for all other parameters.
For comparison with task-specific pre-trained models, we also fine-tuned BART-large555https://huggingface.co/facebook/bart-large on the DialogSum dataset and fine-tuned BARThez666https://huggingface.co/moussaKam/barthez on the DECODA dataset. Details of the training are presented in Appendix A.
Summarization Prompt Design.
| Settings | Prompts |
| WordLimit (WL) [1] | System: Write a summary with not more than X words. User: [Test Dialogue] |
| Human Guideline (HG) for DialogSum | System: Write a summary based on following criteria: the summary should (1) convey the most salient information of the dialogue and; (2) be brief (no longer than 20% of the conversation length) and; (3) preserve important named entities within the conversation and; (4) be written from an observer perspective and; (5) be written in formal language. In addition, you should pay extra attention to the following aspects: 1) Tense Consistency: take the moment that the conversation occurs as the present time, and choose a proper tense to describe events before and after the ongoing conversation. 2) Discourse Relation: If summarized events hold important discourse relations, particularly causal relation, you should preserve the relations if they are also in the summary. 3) Emotion: you should explicitly describe important emotions related to events in the summary. 4) Intent Identification: Rather than merely summarizing the consequences of dialogues, you should also describe speakers’ intents in summaries, if they can be clearly identified. In addition to the above, you should use person tags to refer to different speakers if real names cannot be detected from the conversation. User: [Test Dialogue] |
| Human Guideline (HG) for DECODA | System: Write a conversation-oriented summary in the form of a synopsis expressing both the customer’s and the agent’s points of view, and which should ideally report: 1. The main issues of the conversation: in call center conversations the main issues are the problems why the customer called; their identification constitutes the basis for classifying the call into several different classes of motivations for calling. 2. The sub-issues in the conversation: when in the conversation any sub-issue occurs, it may be there both because it is introduced by the customer or by the agents. 3. The resolution of the call: i.e. if the customer’s problem was solved in that call (first-call resolution) or not. User: [Test Dialogue] |
| Human Guideline with Role (HGR) | The same as the Human Guideline (HG) for DECODA but we changed the begin of instruction to You are an annotator and are asked to write (dialogue summaries) …. |
The different prompts used in our experiments are given in Table 1. The WordLimit (WL) prompt simply constrains the word length of the output to be less than X words [1]. Another prompt we tested was to provide a text only version of the full Human Guideline (HG). For DialogSum, the annotation guideline was directly described in Chen et al. [7]. For DECODA, we contacted the authors [12] and obtained the exact guidelines used to write the synopses. In addition, to examine if giving the system a role helps, we proposed Human Guideline with Role (HGR), which begins with "You are an annotator …". We have also investigated a two-step iterative approach, called HG(R)WL, which first uses the summaries generated by HG or HGR as input, then re-injects them into the model to reduce the length of the output using the WL prompt.
Evaluation Metrics.
We reported the F1 scores of ROUGE-1/2/L [14], which assess the similarity between candidates and references based on the overlap of unigrams, bigrams, and the longest common sequence. We used a publicly available implementation.777https://github.com/google-research/google-research/tree/master/rouge
We also reported the F1 scores of BERTScore [15], which measures the similarity between candidates and references at token level, using BERT’s contextual embeddings.888https://huggingface.co/spaces/evaluate-metric/bertscore For DialogSum, following Chen et al. [13], we used RoBERTa-large [16] as the backbone. For DECODA in French, we used the default multilingual model: bert-base-multilingual-cased [17].
However, few metrics excel across all dimensions, leading recent studies to explore LLMs for evaluating text quality [18]. Following this trend, we adopted an LLMs-as-judge approach, using GPT-4o as the backbone for automatic evaluation on a subset of model predictions.
3 Results
| DialogSum | R1 | R2 | RL | BS | Avg. Len |
| Human Ref. | 53.35 | 26.72 | 50.84 | 92.63 | 21.07* |
| GoodBai | 47.61 | 21.66 | 45.48 | 92.72 | 25.72* |
| UoT | 47.29 | 21.65 | 45.92 | 92.26 | 27.05* |
| BART-large | 47.36 | 21.23 | 44.88 | 91.42 | 27.20 |
| 3.5-WL | 32.92 | 12.45 | 26.66 | 88.78 | 28.69 |
| 3.5-HG | 24.15 | 8.60 | 18.70 | 87.68 | 104.89 |
| 3.5-HGR | 26.43 | 9.27 | 20.50 | 88.16 | 88.85 |
| 3.5-HGWL | 35.10 | 11.52 | 27.39 | 89.72 | 40.02 |
| 3.5-HGRWL | 35.76 | 11.71 | 28.19 | 89.87 | 36.18 |
| 4-WL | 34.94 | 12.38 | 28.99 | 89.15 | 18.63 |
| 4-HGR | 26.25 | 8.01 | 19.86 | 88.44 | 86.78 |
| 4-HGRWL | 35.42 | 9.58 | 28.20 | 90.25 | 19.77 |
| 4o-WL | 35.34 | 11.09 | 28.83 | 89.33 | 18.62 |
| 4o-HGR | 25.53 | 7.92 | 19.34 | 88.21 | 89.67 |
| 4o-HGRWL | 33.80 | 9.18 | 26.96 | 89.83 | 20.23 |
| DECODA | R1 | R2 | RL | BS | Avg. Len |
| Reference | - | - | - | - | 25.40 |
| BARThez | 35.42 | 16.96 | 29.41 | 74.94 | 18.53 |
| 3.5-WL | 33.21 | 12.53 | 24.74 | 72.54 | 37.90 |
| 3.5-HG | 18.42 | 7.18 | 13.68 | 67.72 | 162.83 |
| 3.5-HGR | 16.72 | 6.73 | 12.66 | 66.93 | 190.87 |
| 3.5-HGWL | 31.61 | 11.00 | 23.34 | 71.77 | 46.30 |
| 3.5-HGRWL | 32.74 | 12.22 | 24.09 | 72.30 | 41.76 |
| 4-WL | 35.23 | 13.54 | 27.73 | 73.23 | 23.81 |
| 4-HGR | 20.71 | 8.29 | 15.12 | 69.04 | 136.53 |
| 4-HGRWL | 32.12 | 11.17 | 24.82 | 72.30 | 26.01 |
| 4o-WL | 34.86 | 13.05 | 27.04 | 73.59 | 22.73 |
| 4o-HGR | 19.35 | 7.73 | 13.94 | 68.57 | 160.73 |
| 4o-HGRWL | 31.43 | 10.40 | 23.67 | 72.42 | 24.17 |
3.1 Quantitative Results
DialogSum
The results of the automatic evaluations on the test set are presented in Table 2. The performance of the ChatGPT model with the WordLimit prompt is approximately points lower for R1, points lower for R2, and points lower for RL compared to state-of-the-art pre-trained models. However, BERTScore suggests only a small semantic difference. The results for GPT-4 with the WordLimit prompt are slightly better than those of ChatGPT, except for R2. The results of GPT-4o are comparable to those of GPT-4, with the WL prompt demonstrating superior performance across most metrics, except for BERTScore.
Although the DialogSum guideline states "be brief (no longer than 20% of the conversation length)", the results of the HG prompt indicate that instructions tailored for human annotators may not be suitable when directly used as instructions to generate reference-like summaries with ChatGPT, as responses tend to be longer, leading to lower ROUGE and BERTScore values. This effect could have been reduced by specifying the role of the annotator in the prompt (HGR), which consistently improved performance across all measures.
Regarding the two-step prompt approach HGRWL, which incorporates human annotation instructions as a first step, followed by applying a second prompt to limit word length, the performance aligned with those obtained with a single WordLimit prompt, with even higher R1 and RL scores, suggesting that the GPT model’s word usage is more akin to the references, maintaining the primary content and logical order, albeit with some variability in language usage.
In terms of average length, the WordLimit prompt consistently gave the shortest summaries for both GPT models. GPT-4 and GPT-4o better followed the WordLimit prompt in obtaining a summary of less than words compared to ChatGPT.
DECODA
Results in Table 3 show that the WordLimit prompt produces the best outcomes for both GPT models. However, their performance still falls short of the pre-trained model across all metrics. The difference is less pronounced in the BERTScore than in the ROUGE scores.
When using human guidelines (HG(R)), results diverged from those of the other prompts, resulting in reduced ROUGE and BERTScore. For ChatGPT, HGR showed lower results than HG. The two-step prompting HG(R)WL produced shorter final summaries than HG(R), resulting in higher scores across various measures, although they remain slightly lower than those obtained with the simple WordLimit prompt.
When comparing the GPT models, all three models performed better with the WordLimit prompt, while yielding similar results with the iterative prompt (HGRWL). In terms of average length, the WordLimit prompt consistently gave the shortest summaries for all GPT models. Again, the GPT-4 and GPT-4o models better followed the WordLimit prompt in obtaining a summary of less than words compared to ChatGPT, which always exceeded the limit.
Using LLM-as-judge
Using DeepSeek-R1999https://api-docs.deepseek.com/guides/reasoning_model as our evaluation backbone, we assessed the generated summaries against human summarization guideline criteria (see Appendix B for details and prompts). We evaluated a subset of 20 samples from the DECODA dataset, with results shown in Table 4.
However, we observed that even when the prompt explicitly provided guidelines for judging summaries and specified that only the note should be returned, many evaluation results contained not only scores but also scoring explanations – sometimes in French and sometimes in English.
The GPT models’ generated summaries achieved strong scores across all four criteria. The best results for Faithfulness, Main Issues, and Resolution came from 4-HGRWL, 3.5-WL, and 3.5-HGRWL respectively, while the References performed best on Sub-Issues. Considering average scores across all criteria, 4-HGRWL achieved the highest scores while Barthez obtained the lowest.
| Faithfulness (VN) | Main Issues (VN) | Sub-Issues (VN) | Resolution (VN) | Average (VN) | |
| Reference | 3.65 1.09 (20) | 4.35 1.09 (20) | 3.85 1.27 (20) | 3.15 1.81 (20) | 3.75 1.32 (20) |
| 3.5-WL | 4.00 0.86 (20) | 4.60 0.75 (20) | 3.60 1.10 (20) | 3.35 1.73 (20) | 3.89 1.11 (20) |
| 3.5-HGRWL | 3.20 1.47 (20) | 4.00 1.30 (20) | 3.75 1.21 (20) | 3.95 1.15 (20) | 3.73 1.28 (20) |
| 4-WL | 4.00 1.12 (20) | 4.40 0.82 (20) | 3.42 1.39 (19)† | 2.45 1.70 (20) | 3.57 1.26 (20) |
| 4-HGRWL | 4.15 0.93 (20) | 4.45 0.76 (20) | 3.80 0.83 (20) | 3.50 1.54 (20) | 3.98 1.02 (20) |
| BARThez | 1.70 1.13 (20) | 2.00 1.65 (20) | 1.42 0.84 (19)‡ | 1.60 1.27 (20) | 1.68 1.22 (20) |
†One sample in 4-WL was excluded due to "N/A".
‡One sample in BARThez was excluded due to an error in returning the score.
3.2 Quantitative Results by Variation
| ROUGE-1 | ROUGE-2 | ROUGE-L | BERTScore | |
| Mean (std) | 53.37 (16.75) | 26.73 (20.57) | 44.99 (18.11) | 92.86 (2.68) |
| 4-WL | 34.96 | 12.35 | 28.99 | 89.15 |
| 4-HGR | 26.24 | 8.01 | 19.86 | 88.44 |
| 4-HGRWL | 35.43 | 9.58 | 28.18 | 90.25 |
For the DialogSum dataset, each dialogue is annotated by three annotators, resulting in three reference summaries per dialogue. Despite the shared annotation guideline, human annotators can exhibit variability in writing style. To evaluate whether GPT models follow human summarization guidelines, we compare the models’ scores to the variance observed in the reference summaries for each metric (ROUGE-1/2/L and BERTScore). This approach determines whether the model-generated summaries fall within a reasonable range of human-annotated references, based on the mean and standard deviation of the reference scores.
We evaluated predictions from 4-WL, 4-HGR, and 4-HGRWL, with results presented in Table 5. The first row – Mean (std) – represents the variance among reference summaries from three annotators. Our findings show that 4-WL predictions fall within the reference variance for R2 and RL, 4-HGR only for R2, and 4-HGRWL for R2, RL, and BERTScore. These results demonstrate that, similar to how different annotators exhibit varying styles in writing references, the generated summaries align with the variance distribution of human annotators for certain metrics.
3.3 Manual Error Analysis
We also analyzed the data points where GPT-generated summaries exhibit the greatest discrepancy, characterized by low ROUGE scores but high BERTScore values. For comparison, we selected predictions from two models: 4-WL, which uses a simple WordLimit prompt, and 4-HGRWL, which incorporates human summarization guidelines as an intermediate step. We identified the five examples with the largest discrepancies for DialogSum predicted by 4-WL (Table 6) and 4-HGRWL (Table 15). For DECODA, we present samples predicted by 4-WL (Table 7) and 4-HGRWL (Table 16), both translated from French into English.
| ID | References | Prediction | RL | BS | Diff. |
| 16 | ["#Person2# at first thinks #Person1#’s behaviour cruel but finally joins #Person1#.", "#Person1# plans on playing a trick to others. #Person2# thinks it’s cruel at first but then joins.", "#Person1# is about to make a prank. #Person2# thinks it’s cruel at first but then joins."] | Two friends engage in a potentially offensive prank, causing a disagreement about their choice of fun. | 6.00 | 86.09 | 80.09 |
| 24 | ["#Person1# is interviewing #Person2# about #Person2#’s ability and previous experience.", "#Person1# asks #Person2# about #Person2#’s capacities and past working experience during an interview.", "#Person1# asks #Person2# about #Person2#’s qualification for the job."] | Person 2, an Automobile Designing graduate, has experience in understanding mechanical strength and corrosion resistance. | 4.86 | 83.67 | 78.81 |
| 255 | ["#Person1# and #Person2# are conjecturing who will get sacked based on other employees’ performance and find that it might be themselves. #Person1# is so anxious as #Person1# needs to raise family but #Person2# will be happy to leave.", "#Person1# and #Person2# talk about the company’s upcoming downsizing and discuss who will get sacked. #Person1# is worried but #Person2# doesn’t care.", "#Person1# and #Person2# have rounds of conjectures of whom shall be fired due to the company’s cutback and comes to the conclusion that they both have the biggest possibility. It drives #Person1# crazy but doesn’t worry #Person2#."] | Employees speculate about potential layoffs at their company, considering various colleagues’ relationships with the boss. | 7.15 | 85.90 | 78.76 |
| 48 | [’#Person1# and #Person2# are talking about the low temperature at night, although spring has come.’, ’#Person1# and #Person2# agree that it still felt very cold in spring.’, ’#Person1# and #Person2# talk about the weather and how to keep warm.’] | Despite spring’s arrival, cold nights persist, causing discomfort indoors and outdoors. | 8.02 | 86.53 | 78.50 |
| 67 | [’#Person1# is driving #Person2# to an inn. They talk about their careers, ages, and where they was born.’, ’#Person1# drives #Person2# to an inn and they have a talk. #Person2# is 26 and had a business trip to China. #Person1# is 40 years old American.’, ’#Person1# drives #Person2# from the airport to an inn and they have a casual talk about themselves.’] | A Mexican businesswoman, returning from China, converses with her American-Colombian taxi driver. | 7.22 | 84.96 | 77.74 |
| ID | Reference | Prediction | RL | BS | Diff. |
| 10 | Request for information about a large account order. Transferred to the relevant service. | A seminar organizer is seeking transport cards for their international guests from the RATP. | 0 | 71.13 | 71.13 |
| 110 | Request for information following the procedure after receiving a fine notice. Transferred to the relevant service. | A customer calls customer service to understand why she has to pay five euros after contesting and justifying a bus fine. | 8.33 | 70.86 | 62.53 |
| 107 | request for a refund procedure following a ticket purchase due to the late delivery of the ImagineR pass, with a waiting period of over 3 weeks. Refund possible, contact ImagineR by phone or email. | A customer calls to request a refund for an Orange card purchased while waiting for her Imagine+R card, which arrived late. | 8.00 | 69.35 | 61.35 |
| 32 | T2 circulation | A user inquires about the launch of the T+two line from Porte de Versailles to Val d’Issy. | 8.70 | 69.85 | 61.15 |
| 29 | Request for information on purchasing tickets for school groups. Transferred to the relevant service. | A representative of a departmental association is looking to organize transport for more than 100 classes using public transportation. | 10.26 | 71.05 | 60.80 |
4-WL
First, we look at the summaries generated by 4-WL. We observed in the DialogSum examples with the greatest discrepancies that the references specifically articulate by preserving mentions of #Person1# and #Person2#, summarizing the conversation while retaining named entities. In contrast, GPT-4 predictions often summarize the conversation at a higher, generalized level, using terms like "two friends" or "employees" instead of specific tags. This demonstrates that GPT-generated summaries fail to meet the following guideline: "3) Preserve important named entities within the conversation" when using a simplified prompt.
In addition, GPT-4 predictions don’t match the references concerning the guideline emphasizing "Intent Identification"1010104) Intent Identification: Rather than merely summarizing the consequences of dialogues, you should also describe speakers intents in summaries, if they can be clearly identified. In addition to the above, you should use person tags to refer to different speakers if real names cannot be detected from the conversation. . For instance, in the second example, the prediction references only Person 2, omitting the intent of Person 1, who is interviewing the other individual. In the fourth example, neither Person 1 nor Person 2 is mentioned, and their intents are absent, as the prediction reduces the dialogue to a general remark about the weather. In the fifth example, the prediction fails to follow the "use person tags" rule, summarizing the conversation by identifying the interlocutors as "A Mexican businesswoman" and "her American-Colombian taxi driver". These variations probably influenced the quantitative results evaluated by automated metrics, as evidenced by the high BERTScore values indicating good semantic similarity, while the low ROUGE-L scores suggest minimal overlap in the longest common sequence. This disparity highlights the tendency of GPT-generated summaries using a simple length constraint prompt to capture the gist and key information of the conversation while diverging in lexical and structural alignment with reference summaries.
For the DECODA dataset, the discrepancy may stem from the distinct characteristics of the reference synopses. These summaries focus strictly on issues and resolutions, whereas GPT-generated predictions often adopt a more individual-centered perspective, using terms such as "un organisateur (an organizer)", "une cliente (a client)", and "un représentant (a representative)", etc. This difference results in relatively high BERTScore values, reflecting strong semantic similarity, but low ROUGE-L scores due to minimal word overlap with the references. Additionally, reference synopses are often extremely brief and sometimes incomplete sentences, as seen in the fourth example: "circulation du T2" (T2 circulation). Furthermore, all samples except the fourth ("T2 circulation") fail to conclude the resolution in the generated summaries.
4-HGRWL
Next, we examine the generated summaries of 4-HGRWL. On the DialogSum dataset, we observe better retention of named entities and personal tags (e.g., Person 1) compared to 4-WL, indicating that the intermediate HGR step helps the model to follow specific rules from the human summarization guidelines. However, the issue of overlooking Intent Identification persists, as prediction often focus on one individual’s intent while omitting the other’s, whereas references typically indicate both interlocutors’ intents.
For DECODA, similar to the 4-WL outputs, most generated summaries employ an individual-centered perspective that creates divergence in -gram overlap and semantic similarity. While reference summaries maintain conciseness, model predictions tend toward detail while frequently omitting critical Resolution elements. For example, the first sample omits "Transfer to the relevant department" despite this being an explicit requirement in the summarization guidelines.
In summary, comparing the data points with the greatest discrepancy (low ROUGE-L but high BERTScore) between 4-WL and 4-HGRWL, the latter better adheres to guidelines regarding named entities and personal tags. However, both models share similar shortcomings: in DialogSum, they overlook "Intent Identification"; in DECODA, they use an individual-centered perspective and omit the "Resolution".
3.4 Human Evaluation
Due to the misalignment between automatic measures and human judgment of LLM outputs, as evidenced in news summarization experiments with GPT-3 [20], we set up a human evaluation. To determine if the generated summaries meet predefined instructions, we conducted a human evaluation using criteria derived from the annotation guidelines provided to human annotators for writing reference summaries. We started with the DECODA dataset, as its task-oriented summarization guidelines are more easily adaptable into evaluation criteria. We selected the 10 shortest and 10 longest dialogues, assessing six summaries for each dialogue, including references and predictions from various models: BARThez, 3.5-WL, 3.5-HGRWL, 4-WL, and 4-HGRWL.
Annotation interface
For our human evaluation task, we adapted the POTATO annotation tool [21]. The evaluation interface, shown in Figure 1, features a dialogue at the top, accompanied by six different summaries, including references and model outputs from BARThez, 3.5-WL, 3.5-HGRWL, 4-WL, and 4-HGRWL. To minimize bias, we randomized the order of summaries, assigning them names A, B, C, D, E, and F.
There are four blocks representing the different evaluation criteria, namely Faithfulness, Main issues, Sub-issues and Resolution. In each block, annotators interact with the interface by clicking on each of the six summaries, assigning a score from to to each summary based on its compliance with the specific criterion, being the best.
Three native French speakers took part in the evaluation, all of them graduate students. We presented them with a PDF evaluation guide and explained that the annotations they made would be used for analysis.
Criteria
Annotators assessed each summary based on four criteria: Faithfulness, Main Issues, Sub-Issues, and Resolution. The later three criteria are based on human summarization guidelines, while Faithfulness was added to evaluate the accuracy of the facts presented in relation to the dialogues. Sub-Issues was only evaluated if at least one sub-issue was present in the dialogue.
The evaluation criteria are described below. In the evaluation guide, we also provide detailed explanations for each score of criterion.
Faithfulness: The summary must respect the dialogue at the level of factual information.
Main issues: in call center conversations the main issues are the problems why the customer called; their identification constitutes the basis for classifying the call into several different classes of motivations for calling.
Sub-issues: when in the conversation any sub-issue occurs, it may be there both because it is introduced by the customer or by the agents.
Resolution: i.e. if the customer’s problem was solved in that call (first-call resolution) or not.
Results
| Faithfulness | Main Issues | Sub-issues | Resolution | Overall | |
| Reference | 3.721.12 | 4.151.07 | 3.221.86 | 3.251.60 | 3.681.36 |
| BARThez | 2.171.43 | 2.321.57 | 1.000.0 | 2.221.54 | 2.171.49 |
| 3.5-WL | 4.021.17 | 4.301.18 | 2.671.73 | 4.181.28 | 4.101.28 |
| 3.5-HGRWL | 3.971.15 | 4.470.89 | 3.671.73 | 4.201.13 | 4.191.12 |
| 4-WL | 4.130.96 | 4.370.80 | 1.441.33 | 2.701.79 | 3.621.53 |
| 4-HGRWL | 3.931.02 | 4.300.96 | 2.671.66 | 4.051.27 | 4.031.16 |
Table 8 presents the human evaluation results across four criteria and their overall scores for the 20 evaluated dialogues. A comparison, including all dialogues, the 10 shortest, and the 10 longest dialogues, is available in Table 10 in Appendix. Overall, 3.5-HGRWL achieved the highest score of , though 4-WL slightly excelled in Faithfulness. For the 10 shortest dialogues, 3.5-HGRWL led with an overall score of , while 3.5-WL had the highest overall score at for the 10 longest dialogues.
The Sub-issues criterion is challenging to analyze since sub-issues are not presented in all dialogues, and there was variability in how different evaluators assessed this criterion. Specifically, one annotator identified sub-issues in five dialogues, another in three, and a third in one dialogue. In addition, the annotators generally preferred the output from GPT models over the reference summaries, with BARThez being less favored overall. However, it was observed that 4-WL received the lowest scores for the Resolution criterion specifically in the 10 longest dialogues.
The annotators’ preference for GPT-generated summaries may stem from their use of more natural and fluid language. In the DECODA dataset, the reference synopses exhibit highly specific and concise characteristics. For example, as shown in Table 14, reference synopses are short and to the point: "Request for train timetables from Maux station to Gare de l’Est station at a given time” (Demande d’horaires de train de la gare de Maux à la gare de l’Est à une heure donnée). In some cases, the synopses are not even complete sentences. This contrast in style could explain the annotators’ tendency to favor the more natural and cohesive language of GPT-generated outputs.
Inter-Annotator Agreement (IAA)
We calculated inter-annotator agreement (IAA) scores across all three annotators, with results shown in Table 9. The analysis reveals better agreement on the Resolution criterion, while there is considerable variation in their annotations for Sub-issues.
| A1-A2 | A1-A3 | A2-A3 | |
| Faithfulness | 0.422 | 0.273 | 0.315 |
| Main issues | 0.363 | 0.290 | 0.398 |
| Sub-issues | 0.140 | 0.445 | -0.175 |
| Resolution | 0.546 | 0.503 | 0.463 |
4 Discussion and Conclusion
Human evaluation results show that GPT models (ChatGPT, GPT-4 and GPT-4o) can follow human summarization guidelines to some extent. Their summaries were preferred over task-specific pre-trained models and even outperformed reference summaries. The longer outputs of GPT models, compared to BARThez, likely contribute to more comprehensive coverage of issues and resolutions, enhancing faithfulness to dialogues. However, GPT models underperform task-specific pre-trained models on both ROUGE and BERTScores metrics. This discrepancy between automatic metrics and human judgments aligns with previous findings, reinforcing (1) the continued necessity of human evaluation for text generation tasks, and (2) the critical need for developing better-aligned automatic evaluation metrics.
Additionally, GPT models demonstrate proficiency in following human guidelines. In the overall evaluation, using human guidelines – HGRWL tends to produce better results compared to the simpler WordLimit approach. Interestingly, GPT-4 does not consistently outperform ChatGPT; 4-WL only surpasses 3.5-WL in the 10 shortest dialogues. Nevertheless, GPT-4 adheres better to word length constraints specified in the instructions.
Our findings also highlight the inherent subjectivity in assessing summary quality, as human evaluators often preferred GPT-generated summaries over reference texts – a preference partially attributable to stylistic differences.
By analyzing data points where GPT-generated summaries (4-WL and 4-HGRWL) exhibit significant discrepancies – characterized by low ROUGE scores but high BERTScore values – we found that the models effectively summarize key dialogue information and are semantically similar to the references. However, they sometimes neglect specific rules: in DialogSum, they miss named entities, intent identification, and personal tags; in DECODA, some summaries focus on individual locutors’ perspectives and fail to present resolutions. The intermediate step of using HGR helped the model better adhere to the specific rules of human summarization guidelines, such as the use of named entities and personal tags in DialogSum.
In future work, we aim to explore LLM-based automatic evaluation measures and develop methods to ensure stable and consistent responses from them, better aligned with human judgment. It would be possible to explore the use of LLM-as-judge approaches as complementary metrics to enhance the robustness of automatic metrics. Additionally, it would also be interesting to investigate LLM behavior in following different summarization guidelines, such as those involving lexical perturbations, to better interpret and explain their decision-making processes.
Limitations
In terms of model selection, our analysis was limited to GPT models: ChatGPT, GPT-4 and GPT-4o. Future studies could benefit from a broader exploration of LLMs, encompassing models with varying parameter sizes and accessibility.
Ethics Statement
The DialogSum corpus we used in this study is composed of resources freely available online without copyright constraint for academic use. The training and validation sets of the DECODA dataset can be downloaded from their website, on acceptance of the corresponding usage and sharing terms, and the test set is available only on request from the authors.
We chose these two corpora because they have different communication objectives, are produced in different languages and belong to different domains. In addition, they were both used in challenges, making it easier to compare results.
Appendix A Experiment details: BART-based models
DialogSum
We fine-tuned the BART-Large model111111https://huggingface.co/facebook/bart-large on the DialogSum dataset with the following hyperparameter settings: a learning rate of , 15 epochs, a training batch size of 2, and evaluation/test batch size of 8. The maximum source length was set to 1024 tokens, and the maximum target length to 128 tokens, with a default length penalty value of 1.0. The experiments were conducted on the NVIDIA Quadro RTX 6000 GPU and took approximately 2.5 hours per run.
DECODA
We fine-tuned the BARThez121212https://huggingface.co/moussaKam/barthez model available in Huggingface on the preprocessed DECODA dataset. Experiments were run on 1 NVIDIA Quadro RTX 8000 48Go GPU server. We used the default parameters for summarization tasks, which are: initial learning rate of , with a train batch size and eval batch size of 4, and a seed of 42. We used Adam optimizer with a linear learning scheduler. In details, we fine-tuned BARThez (base architecture, 6 encoder and 6 decoder layers) for 10 epochs and saved the best one having the lowest loss on the dev set. Each training took approximately 25 minutes.
Appendix B LLM-as-judge
For the LLM-as-judge approach, we adapted the G-Eval framework [18], modifying its criteria to align with DECODA’s human summarization guidelines (Main Issues, Sub-Issues, Resolution) while additionally incorporating Faithfulness. Each criterion was evaluated using distinct prompts.
As an example, Figure 2 shows the specific prompt used to assess the Main Issues dimension.
You will receive a dialogue. You will then receive a written summary of this dialogue.
Your task is to evaluate this summary against a single criterion.
Please read these instructions carefully and ensure you understand them. Keep this document open during evaluation and refer to it as needed.
Evaluation Criterion:
Main issues (1-5) - The summary should capture the main topics of the conversation. These are the problems for which the customer called. Identifying these issues forms the basis for classifying the call into different motivation categories. The main problem of a call should be prioritized for inclusion in the summary.
The 5-point Likert scale is as follows:
1.
Incorrect or missing: The main issues of the call are presented incorrectly, are missing, or are completely different from those discussed during the call.
2.
Confusing: The main issues of the call are presented in a confusing or ambiguous way, making it difficult to understand the problems addressed.
3.
Correct but incomplete: The main issues are correctly identified, but the presentation lacks details or context needed for full understanding.
4.
Correct but could be improved: The main issues are correctly identified and clearly presented, but some details or clarifications are missing for complete understanding.
5.
Precise and concise: The main issues are presented clearly, concisely, and precisely, enabling full understanding of the problems addressed during the call.
Example:
Source text:
{{Document}}
Summary:
{{Summary}}
Evaluation form (ONLY ratings):
Main issues:
Appendix C Human Evaluation
C.1 Full results
The full human evaluation results for the DECODA dataset are detailed in Table 10, categorized into three groups: all 20 dialogues, the 10 shortest dialogues, and the 10 longest dialogues.
| Faithfulness | Main Issues | Sub-issues | Resolution | Overall | |
| All Dialogues | |||||
| Reference | 3.721.12 | 4.151.07 | 3.221.86 | 3.251.60 | 3.681.36 |
| BARThez | 2.171.43 | 2.321.57 | 1.000.0 | 2.221.54 | 2.171.49 |
| 3.5-WL | 4.021.17 | 4.301.18 | 2.671.73 | 4.181.28 | 4.101.28 |
| 3.5-HGRWL | 3.971.15 | 4.470.89 | 3.671.73 | 4.201.13 | 4.191.12 |
| 4-WL | 4.130.96 | 4.370.80 | 1.441.33 | 2.701.79 | 3.621.53 |
| 4-HGRWL | 3.931.02 | 4.300.96 | 2.671.66 | 4.051.27 | 4.031.16 |
| 10 Short Dialogues | |||||
| Reference | 3.531.07 | 4.071.11 | 5.00 | 3.171.70 | 3.601.37 |
| BARThez | 2.271.39 | 2.201.63 | 1.00 | 2.271.57 | 2.231.51 |
| 3.5-WL | 3.831.32 | 4.131.33 | 1.00 | 4.471.04 | 4.111.29 |
| 3.5-HGRWL | 4.271.01 | 4.670.80 | 5.00 | 4.371.10 | 4.440.98 |
| 4-WL | 4.370.89 | 4.470.78 | 5.00 | 4.171.21 | 4.340.97 |
| 4-HGRWL | 4.100.92 | 4.630.61 | 5.00 | 4.201.30 | 4.321.00 |
| 10 Longest Dialogues | |||||
| Reference | 3.901.16 | 4.231.04 | 3.001.85 | 3.331.52 | 3.761.36 |
| BARThez | 2.071.48 | 2.431.52 | 1.000.0 | 2.171.53 | 2.121.48 |
| 3.5-WL | 4.201.00 | 4.471.01 | 2.881.73 | 3.901.45 | 4.081.27 |
| 3.5-HGRWL | 3.671.21 | 4.270.94 | 3.501.77 | 4.031.16 | 3.951.19 |
| 4-WL | 3.900.99 | 4.270.83 | 1.000.0 | 1.230.77 | 2.961.65 |
| 4-HGRWL | 3.771.10 | 3.971.13 | 2.381.51 | 3.901.24 | 3.761.24 |
C.2 Radar chart
The results for the 20 selected samples are presented in a radar chart in Figure 3. This chart displays the evaluation of six summaries (including both reference and those generated by different systems) on four criteria as well as overall scores. The visual representation indicates that GPT-generated summaries are generally preferred by human annotators compared to the task-specific pre-trained model, BARThez, and that they perform better than reference summaries on some criteria, such as Faithfulness.
Appendix D Example Analysis of Outputs
DialogSum
We present three examples in the following tables. We observe that, when using prompts designated as HG(R), both GPT models tend to generate long summaries that are too precise and detailed, sometimes exceeding the length of the original conversation. Yet the instructions explicitly state that the summary should not exceed 20% of the original dialogue. Consequently, GPT models demonstrate an inability to adhere to and follow the prescribed human summarization guidelines.
In detail, Table 11 reveals that the summary generated by ChatGPT with the WordLimit prompt is of higher quality than that produced by GPT-4. It talks not only about the benefits of the job, but also about its demanding schedule. However, for ChatGPT, we found HGWL to be quite good, whereas HGRWL seems to tell the story from Person1’s point of view instead of Person2’s. As for GPT-4, the summary predicted by HGRWL isn’t very good, as it only talks about the benefits, omitting the demanding schedule of this new job.
In Table 12, the predictions of WordLimit prompt are good for both GPT models, while the other generated summaries are too long; in particular, this conversation is really short. In Table 13, the predictions of WordLimit prompt are good for both GPT models, and those of the prompt HG(R)WL for both GPT models also seem good.
| test_20 |
| Dialogue: #Person1#: Hey, Frank. I heard you got a new job. #Person2#: Yeah, Judy. I will be working for the Post Office. It’s not a bad job. #Person1#: Is it true that you have a heavy work schedule? #Person2#: That’s right. I am supposed to work at 5am everyday, and I only get 45 minutes for lunch. #Person1#: So, why did you apply for such a demanding job? #Person2#: Well, the government offers its employees excellent health insurance benefits. #Person1#: Oh, I see. And can your family members utilize the health insurance, too?#Person2#: Yeah, that’s the best part. All of my children can get free medical care. #Person1#: That’s a great employment benefit! #Person2#: Now you know why I wanted to work for the Post Office! |
| Summary1: Frank got a new job and is telling Judy not only the heavy schedule but also the benefits of this job. |
| Summary2: Frank explains to Judy the reason why he took a job at the Post office is that the government offers excellent health insurance benefits for its employees. |
| Summary3: Frank describes his new job’s schedule and insurance at the Post Office to Judy. |
| Generated Summaries – ChatGPT |
| 3.5-WL: Person2 got a new job at the Post Office with a demanding work schedule, but the excellent health insurance benefits for family members made it worth it. |
[…]
| 3.5-HGWL: Person2 tells Person1 about their new job at the Post Office, mentioning the demanding work schedule but highlighting the excellent health insurance benefits for their family. |
[…]
| Generated Summaries – GPT-4 |
| 4-WL: Frank got a new job at the Post Office for its excellent health insurance benefits for his family. |
[…]
| 4-HGRWL: Frank accepts a demanding Post Office job for its excellent health insurance benefits, covering his children’s medical care. |
| test_151 |
| Dialogue: #Person1#: Have you enjoyed your weekend? #Person2#: Yes, it’s been marvelous. It really was very kind of you to invite me. I hope it hasn’t been too much trouble. #Person1#: Not at all. We’ve really enjoyed having you. I hope you’ll come and stay again next time you’re in Changsha. |
| Summary 1: #Person2# tells #Person1# #Person2# enjoyed the weekend organized by #Person1#. |
| Summary 2: #Person1# and #Person2# has enjoyed a nice weekend together. |
| Summary 3: #Person1# invited #Person2# to spend the weekend and #Person2# has enjoyed it. |
| Generated Summaries – ChatGPT |
| 3.5-WL: Person2 had a marvelous weekend and thanked Person1 for the invitation, hoping to stay again in the future. |
[…]
| 3.5-HGWL: Person2 had a marvelous weekend and thanked Person1 for the invitation. Person1 reassured them and invited them to visit again. |
[…]
| Generated Summaries – GPT-4 |
| 4-WL: Person1 invites Person2 for a weekend stay in Changsha, which Person2 thoroughly enjoys. |
[…]
| 4-HGRWL: Person1 invites Person2 again after a pleasant weekend visit in Changsha. |
| test_440 |
| Dialogue: #Person1#: Man, I’m freaking out! You gotta help me! #Person2#: Whoa, whoa, take it easy, relax. Geez, you’re sweating like a pig! What’s going on? #Person1#: I can’t go through with this! I just can’t! I’m not ready for marriage! What was I thinking? I’m only thirty five years old! I’ve got my entire life ahead of me, adventures waiting! I can’t settle down yet! #Person2#: What are you talking about? It wasn’t more than a month ago that you were rambling on about how you are tired of living the life of a bachelor and how you envy your friends that have a family! #Person1#: I know. I know! #Person2#: Let’s think this through. First of all, you cannot leave Amy at the altar. Not only will she be humiliated and kill you, but she is the woman that you love and the woman of your dreams! Why would you want to end that or jeopardize it like this? Second of all, you are just getting cold feet. You know deep down inside that you want to marry her, so cut the crap and do it! #Person1#: You’re right. I’m being crazy! Ok, I’m ready, let’s do this! #Person2#: Great! Phew! That was a close one! You had me going there for a minute I thought I was gonna have to slap some sense into you. |
| Summary 1: #Person1# is not ready for marriage. #Person2# reminds him of what he said a month ago and what will happen if he leaves Amy. After listening to #Person2#, #Person1# is ready for marriage. |
| Summary 2: #Person1# feels anxious about future marriage and sweats a lot. #Person2# helps #Person1# to think through the current situation and get ready for the marriage. |
| Summary 3: #Person1# is freaking out before his marriage with Amy and #Person2# helps him get over it by helping him think through all the details. |
| Generated Summaries – ChatGPT |
| 3.5-WL: Person 1 is having second thoughts about getting married, but Person 2 convinces them to go through with it. |
[…]
| 3.5-HGWL: Person1 is anxious about getting married but Person2 reassures them and reminds them of their love for their partner. |
[…]
| Generated Summaries – GPT-4 |
| 4-WL: A man panics before his wedding, but his friend calms him down and convinces him to go through with it. |
[…]
| 4-HGRWL: Person1, fearing marriage, is reassured by Person2 and decides to proceed with his wedding to Amy. |
DECODA
We present an example of the sample FR_20091112_RATP_SCD_0826 in Table 14. We note that the synopses generated with WordLimit and HG(R)WL prompts for both GPT models are good. The synopses generated with HG(R) prompt follow the structure of the given instructions: first talk about the main issue, then the sub-issues of the conversation, and finally the resolution of the call, even the predictions are long, but make the two-step prompt approach – HG(R)WL generating quite good synopses.
| Dialogue: Spk_A bonjour |
| Spk_B oui bonjour madame |
| Spk_B je vous appelle pour avoir des horaires de train en+fait c’ est pas pour le métro je sais pas si vous pouvez me les donner ou pas |
| Spk_A trains SNCF |
| Spk_B oui trains SNCF oui |
| Spk_A vous prenez quelle ligne monsieur |
| Spk_B euh la ligne euh enfin en+fait c’ est pas SNCF enfin c’ est Île-de-France quoi je sais pas comment ils appellent ça |
| Spk_B RER voilà c’ est pour les RER |
| Spk_B voilà et euh je prends la ligne euh Meaux de Meaux pour aller à Paris je sais plus c’ est laquelle c’ est la |
| Spk_B E je crois |
| Spk_A d’accord et vous voulez donc les horaires euh pour quel jour et à quel moment monsieur |
| Spk_B euh là pour euh tout à l’ heure euh pour euh aux environs de dix-sept heures en partant de la gare de Meaux |
| Spk_B euh vers la Gare+de+l’Est à Paris |
| Spk_A alors vous partez de Meaux et vous allez donc à la Gare+de+l’Est Spk_A et vous voudriez les horaires Spk_B voilà |
| Spk_A vers euh dix-sept heures Spk_A alors moi je peux regarder ce+que Spk_B ouais dix-sept heures ouais |
| Spk_A j’ ai comme horaires un instant monsieur Spk_A s’il+vous+plaît Spk_B d’accord il y a |
| Spk_B pas de souci |
| Spk_A monsieur s’il+vous+plaît |
| Spk_B oui |
| Spk_A donc voilà ce+que j’ ai comme horaires moi vous avez donc un départ à la gare de Meaux donc à seize heures quarante-neuf |
| Spk_B seize heures quarante-neuf d’accord |
| Spk_A et après vous avez celui de dix-sept heures dix-neuf |
| Spk_B alors seize heures quarante-neuf dix-sept heures dix-neuf Spk_A OK d’accord Spk_B oui |
| Spk_B ben je vous remercie |
| Spk_A mais je vous en prie Spk_B bonne journée madame |
| Spk_A merci à vous aussi monsieur au+revoir |
| Spk_B au+revoir |
| synopsis 1: Demande d’horaires de train de la gare de Maux à la gare de l’Est à une heure donnée. |
| synopsis 2: horaires RER E de Meaux à la Gare de l’Est |
| Generated Summaries – ChatGPT |
| 3.5-WL: Un homme appelle pour demander les horaires des trains RER entre Meaux et Paris. Il obtient les horaires pour un départ à 16h49 et à 17h19. |
[…]
| Le problème principal de l’appel est la demande des horaires de train pour la ligne RER de Meaux à Paris. Les sous-problèmes incluent la clarification de la ligne de train (RER) et la spécification de l’heure de départ (aux environs de 17 heures). |
| L’appel est résolu au cours de cet appel, car l’agent fournit les horaires demandés au client. |
| 3.5-HGWL: Le client appelle pour obtenir les horaires de train de la ligne RER de Meaux à Paris. L’agent lui propose deux options de départ : 16h49 ou 17h19. L’appel est résolu. |
[…]
| Generated Summaries – GPT-4 |
| 4-WL: Un homme appelle pour connaître les horaires de train de la ligne RER E entre Meaux et Paris pour le jour même. |
[…]
| 4-HGRWL: Un client a demandé les horaires des trains RER E de Meaux à Paris, l’agent a fourni les informations nécessaires. |
Appendix E Output Length
In Figure 4 and Figure 5, we present statistics on the number of words in summaries generated using different prompts on the DialogSum and DECODA datasets compared to reference summaries, respectively.
The summaries generated by ChatGPT are all longer than the reference summaries. The use of WordLimit prompt showed a shorter average output length, and its use as a second step makes the summaries generated with HG(R) more synthetic (HGRWL).
As for GPT-4, it better respects the instructions regarding output length, we can see that the distribution of 4-WL and 4-HGRWL is more restricted in both figures, while 4-HGR exceeds the length of the reference summaries.
Appendix F Manual Error Analysis
F.1 DialogSum Discrepancy Data Points
Table 15 presents the DialogSum cases where GPT-generated summaries exhibit the most significant scoring discrepancies (low ROUGE-L but high BERTScore values) for 4-HGRWL.
| ID | References | Prediction | RL | BS | Diff. |
| 246 | ["Tim"s project is about how to lead a greener life. #Person1# consults about his motivation and his findings at school. Tim finds there are many aspects that can be improved for his school.", "Tims is talking about his greener life. He now rides his bike to school and recommends teachers to send electrical paper rather than paper ones. And he thinks the school shouldn"t have machines with bottled drinks.", "Tim chooses a green-life-related project because he wants to see what he could do in his own life. He"s started being greener with himself and influences others."] | #Person2# proposes eco-friendly changes at school and personal life, inspired by environmental science lessons. | 6.77 | 86.81 | 80.04 |
| 232 | ["#Person1# asks #Person2# about past experiences of living in different cities around the world.", "#Person1# is asking #Person2# about the places that #Person2# has stayed or lived.", "#Person1# and #Person2# talk about #Person2#"s experience. #Person2# moves a lot. #Person1# finds it interesting."] | #Person2# shares their life journey from Chicago to Los Angeles and Munich with #Person1#. | 9.29 | 89.26 | 79.96 |
| 283 | ["#Person2# wants to buy some clothes but #Person1#"s recommendation doesn"t interest #Person2#.", "#Person1# gives #Person2# a few suggestions on clothes but none is taken.", "#Person2# is not happy with #Person1#"s recommendeds."] | Salesperson suggests brown, silk, and cotton dresses to a customer who prefers a white dress for the warm season. | 4.17 | 83.96 | 79.79 |
| 433 | ["Amy is talking about her first job experience and later job with #Person1#.", "Amy tells #Person1# about her first job experience and her current job.", "Amy shares her former work experience and her current job with #Person1#."] | Amy, after starting as a secretary in 1998, became a Personnel Manager, then moved to Sales due to salary dissatisfaction. | 6.19 | 85.7 | 79.51 |
| 379 | ["#Person1# interviews Vet, a high school girl who has become a mother. She shows her love to her daughter and talks about her future plans and gives advice to teens.", "Vet tells her own experience as a young mother and how her daughter changes her. Then she tells #Person1# her future goals and dreams. And she also gives teens advice.", "#Person1# is interviewing Vet and asking her about the ways she deals with stress as a young mother, feelings of having a baby, plans for the future, her favorite quote, and advice for teens."] | #Person2# discusses stress coping mechanisms, the impact of motherhood, future ambitions, and advises teenagers on self-love and independence. | 11.83 | 89.75 | 77.92 |
F.2 DECODA Discrepancy Data Points
Table 16 displays English translations of DECODA cases where GPT-generated summaries exhibit the most significant scoring discrepancies (low ROUGE-L but high BERTScore values) for 4-HGRWL.
| ID | Reference | Prediction | RL | BS | Diff. |
| 10 | Inquiry about a corporate account order. Transfer to the relevant department. | A representative from the Ministry of Foreign Affairs is organizing a seminar in Paris and contacts RATP to obtain transport cards for participants. | 0.00 | 70.14 | 70.14 |
| 135 | Route request from Drancy station to Gare du Nord. Clarification of explanations following a previous call. | A customer confused about RER B connection to reach Drancy is reassured they can travel directly from Bourg-la-Reine. | 4.55 | 70.29 | 65.75 |
| 29 | Inquiry about purchasing tickets for school groups. Transfer to the relevant department. | A school association representative calls to understand public transport procedures for students. The agent proposes a solution and transfers the call. | 8.51 | 72.83 | 64.32 |
| 68 | Wrong number - caller meant to dial social security at 36 46 but misdialed | A customer mistakes RATP customer service for Social Security regarding a health card issue in the Antilles. | 4.65 | 66.03 | 61.38 |
| 140 | No ImaginR card subscription by mail, only at ticket counters in all stations | A customer requested an Imagine+R application by mail. The agent explained it’s impossible but provided the service number for more information. | 4.17 | 65.54 | 61.38 |
Tables 17 and 18 present the original French DECODA cases where GPT-generated summaries exhibit the most significant scoring discrepancies (low ROUGE-L but high BERTScore values) for 4-WL and 4-HGRWL respectively, as analyzed in Section 3.3.
| ID | Reference | Prediction | RL | BS | Diff. |
| 10 | Demande d’information sur une commande grands comptes. Transfert vers le service concerné. | Un organisateur de séminaire cherche à obtenir des cartes de transport pour ses invités internationaux auprès de la RATP. | 0 | 71.13 | 71.13 |
| 110 | Demande de renseignements suite à la procédure suivi après la réception d’un courrier d’amende. Transfert vers le service concerné. | Une cliente appelle le service clientèle pour comprendre pourquoi elle doit payer cinq euros suite à une amende de bus contestée et justifiée. | 8.33 | 70.86 | 62.53 |
| 107 | demande procédure pour remboursement suite à achat tickets car envoi tardif du passe ImaginR, attente au-delà de 3 semaines donc possible, appeler ImaginR ou par mail | Une cliente appelle pour demander le remboursement d’une carte Orange achetée en attendant sa carte Imagine+R, arrivée en retard. | 8.00 | 69.35 | 61.35 |
| 32 | circulation du T2 | Un utilisateur se renseigne sur la mise en service du T+deux de la Porte de Versailles à Val d’Issy. | 8.70 | 69.85 | 61.15 |
| 29 | Demande de renseignements sur l’achat de ticket pour des groupes scolaires. Transfert vers le service concerné. | Un représentant d’une association départementale cherche à organiser le transport de plus de 100 classes via les transports en commun. | 10.26 | 71.05 | 60.80 |
| ID | Reference | Prediction | RL | BS | Diff. |
| 10 | Demande d’information sur une commande grands comptes. Transfert vers le service concerné. | Un représentant du Ministère des Affaires étrangères organise un séminaire à Paris et contacte la RATP pour obtenir des cartes de transport pour les participants. | 0 | 70.14 | 70.14 |
| 135 | Demande d’itinéraire gare de Drancy gare du Nord. Précision des explications suite à un précédent appel. | Un client confus sur l’interconnexion du RER B pour se rendre à Drancy est rassuré qu’il peut voyager directement depuis Bourg-la-Reine. | 4.55 | 70.29 | 65.75 |
| 29 | Demande de renseignements sur l’achat de ticket pour des groupes scolaires. Transfert vers le service concerné. | Un représentant d’une association scolaire appelle pour comprendre la procédure de transport en commun pour les étudiants. L’agent propose une solution et transfère l’appel. | 8.51 | 72.83 | 64.32 |
| 68 | appel erroné l’appelant ayant voulu joindre l’assurance maladie au 36 46 et ayant mal tapé le numéro | Un client confond le service client de la RATP avec la Sécurité Sociale, concernant un problème de carte vitale aux Antilles. | 4.65 | 66.03 | 61.38 |
| 140 | aucun abonnement de la carte ImaginR par voie postale, seulement au guichet dans toutes les gares et stations | Un client a demandé un dossier Imagine+R par courrier. L’agent a expliqué que c’est impossible mais a fourni le numéro du service pour plus d’informations. | 4.17 | 65.54 | 61.38 |
Appendix G Datasets: License or Terms of Use
The DialogSum corpus we used in this study is composed of resources freely available online without copyright constraint for academic use. The training and validation sets of the DECODA dataset can be downloaded from their website131313https://pageperso.lis-lab.fr/benoit.favre/cccs/, upon acceptance of the corresponding usage and sharing terms, and the test set is available only on request from the authors.